Decide whether or not each equation has a circle as its graph. If it does give the center and the radius.
x^2+8x+y^2-6y+16=0
acrown1@umbc.edu

Answer by fractalier(6550)

Okay from
x^2 + 8x + y^2 - 6y + 16 = 0
we complete the squares twice and get
x^2 + 8x + 16 + y^2 - 6y + 9 + 16 = 0 + 16 + 9
now simplify and form squares and get
(x + 4)^2 + (y - 3)^2 = 9
So this is a circle with center at (-4, 3) and radius 3.